When both arms were locked, we found that ITMY optical lever was very off-center. This seems to have happened after the c1susaux rebooting we did in June 17th. I opened the ITMY table and realigned the OPLev beam to the center when the arm was locked. I repeated this process for BS, PRM and ETMY. I did PRM because I've known that we have been keeping its OpLev off. The reason was clear once I opened the table. The oplev reflection beam was hitting the PD box instead of the PD. After correcting, I was able to swithc on PRM opLev loops and saw normal functioning.